# Strong Bow Shocks: Turbulence and An Exact Self-Similar Asymptotic

### v2

@@ 1 Introduction
Astrophysical bow shocks are ubiquitous; they form at stars flying supersonically through the interstellar medium (e.g., van Buren & McCray, 1988; Kobulnicky et al., 2010; Mackey et al., 2012, 2015), around planets (see e.g., Tsurutani & Stone, 1985; Treumann & Jaroschek, 2008, for comprehensive reviews), or colliding wind binaries (e.g., Stevens et al., 1992; Myasnikov et al., 1998; Gayley, 2009; Sasaki et al., 2023), to name a few scenarios.
@@
The bow shock with infinite Mach number is a classic problem dating back many decades with analytical formulations discussed in the works of Chester (1956a, b), Freeman (1956),van Dyke (1957), Hayes & Probstein (1959), and Yakura (1962), to name a few of the earliest works, and an asymptotic steady-state solution in the astrophysical context was explored with direct hydrodynamical simulation in Yalinewich & Sari (2016, hereafter YS16). To find the shape of the bow shock, YS16 propose the thought experiment: replace an object moving through a cold medium by a properly timed series of explosions. Assuming that the resulting blast waves are sandwiched between each other, YS16 get the bow shock shape $z\propto r^{2}$. YS16 was an important motivation for us, although formally speaking we do not use any of their results. Unlike YS16, we derive an exact self-similar asymptotic for the strong bow shock, and we also discuss turbulence that necessarily manifests near the obstacle head.
@@ 2 Theory
One would think that a mathematical description of a bow shock created by a stellar wind is insurmountably difficult. There are at least three surfaces of discontinuity in the flow: two shocks and one tangential discontinuity  [^1] The terminology and notations of this section follow Landau & Lifshitz (1959)., see Figure 1. The shapes of these singular surfaces are determined by the flow in between them, and the flow itself depends on the shapes of the singular surfaces. “Mathematically such problems are altogether inaccessible to our present analytical techniques” wrote von Neumann, before presenting his exact solution of the strong explosion problem (von Neumann, 1947).
@@ 2.1 Turbulence
Assume that the flow is laminar. Then, in the frame of the star, the flow is steady. In this frame, the cold stellar wind collides with the cold interstellar medium (ISM) wind, Figure 1. Before colliding, the two winds shock and pressurize. The pressurized ISM and the pressurized stellar gas touch each other along a steady tangential discontinuity surface. The gas velocity normal to the tangential discontinuity is zero, the pressure doesn’t jump across the discontinuity, but the tangential velocity does jump. In three space dimensions, even a finite-Mach tangential discontinuity is unstable. According to (Landau & Lifshitz, 1959, section 84, problem 1), this was shown by Syrovatskii (1954). In an ideal inviscid gas, the instability of the tangential discontinuity manifests as turbulence. The instability and the resulting turbulence are essentially 3D; we will need a 3D simulation to confirm turbulence in Section 3.
@@ 2.2 Asymptotically Self-Similar Bow Shocks
where $\gamma$ is the adiabatic index.
@@
As written, Equations (1) – (3) with the boundary conditions (4) – (7) have no meaningful solutions, because the obstacle creating the bow shock is not specified. We will show, however, that at $z\rightarrow+\infty$ a self-similar non-singular asymptotic solution does exist. Namely, for $z\rightarrow+\infty$, one can use the following expansion in small $R^{\prime}$:
@@
This is because, to leading order, $v=1$ at the shock upstream and, to leading order, the Bernoulli equation gives $v={\rm const}$ along a streamline. Thus, asymptotically the bow shock slows down to the ISM velocity. This doesn’t mean that the bow shock becomes “invisible” in the far downstream. Our bow shock is assumed strong, so the maximal density at any fixed arbitrarily large $z$ is still a factor of a few (4 for $\gamma=5/3$) greater than the unperturbed ISM density.
Figure 2: The full structure of the bow show for a rigid obstacle plotted at $t=1,000$. The bow shock wings extend to a maximum of 20 characteristic radii from the symmetry axis.
@@
Equations (15) – (20) are isomorphic to a cylindrically symmetric blast wave; $z$ plays the role of time. An exact solution can immediately be deduced, analogous to the Sedov-von Neumann-Taylor solution (Landau & Lifshitz (1959), section 106):
@@
We show in the Appendix that the focal length of the parabola, $f$, is given by a certain combination of the bow shock parameters. For
@@
Figure 3: The full structure of the bow show for an obstacle with $v_{\rm wind}=0.1$ plotted at $t=1,000$. The bow shock wings are wider than those of the rigid bow shown in Figure 2 due to inner-outer wind interaction.
@@ 3.2 Results
In Figures 2 and 3 we plot the two edge cases considered in this work. Namely, the full structure of the bow shock for the rigid body and the windiest body, i.e, $v_{\rm wind}=0.1$, respectively. In Figure 2 the density field behind the rigid body is evacuated (i.e., $\rho\ll 1$), a testament to the robustness of the immersed boundary method invoked to model the boundary layer. In Figure 3, the density field only gradually decreases to values $\rho\lesssim 0.1$ at large distances $z\leq-150$. This hints at the fact that the windy obstacles approach the asymptotic solutions slower than the perfectly rigid obstacle since the downstream is polluted by turbulent streams of ablated material.
@@ 4 Discussion
This work has shown that the far asymptotic structure of the bow shock with infinite Mach number is self-similar and can be described analytically. After derivation of the exact asymptotic solution, we confirm the theory with high-resolution hydrodynamical simulations for a rigid obstacle and a windy obstacle. The rigid obstacle shows near perfect agreement with the theory with a maximal percent difference of 7 %. The windy obstacle shows rough agreement with the theory, but this is to be expected since the presence of an inner wind introduces turbulence.
@@
Having constructed and verified this theory expounding the asymptotic shape of non-magnetic astrophysical bow shocks, the astrophysical implications can be quite expansive. For example, a current theory explaining the sources behind quasi-periodic eruptions (QPEs; Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021; Chakraborty et al., 2021) are stars that plunge into active galactic nuclei (AGN; e.g., Lu & Quataert, 2023; Linial & Metzger, 2023; Tagawa & Haiman, 2023). The supersonic star creates a bow shock whose thermal energy is radiation dominated and this model is used to explain the radiative properties of some QPEs. In the future, we plan to utilize our theory to explore the fundamental problem of stars plunging through AGN disks in order to glean some of the physics of the stellar entry and exit. This should provide constraints on the á la mode QPE models proposed.
The authors thank Tamar Faran, Yuri Levin, and Zoltan Haiman for useful discussions. MD acknowledges a James Author Fellowship from NYU’s Center for Cosmology and Particle Physics, and thanks the LSST-DA Data Science Fellowship Program, which is funded by LSST-DA, the Brinson Foundation, and the Moore Foundation; his participation in the program has benefited this work.

### v3

@@ 1 Introduction
Astrophysical bow shocks are ubiquitous; they form at stars flying supersonically through the interstellar medium (e.g., van Buren & McCray, 1988; Kobulnicky et al., 2010; Mackey et al., 2012, 2015), around planets (see e.g., Tsurutani & Stone, 1985; Treumann & Jaroschek, 2008, for comprehensive reviews), or colliding wind binaries (e.g., Stevens et al., 1992; Myasnikov et al., 1998; Gayley, 2009; Sasaki et al., 2024), to name a few scenarios.
@@
The bow shock with infinite Mach number is a classic problem dating back many decades with analytical formulations discussed in the works of Chester (1956a, b), Freeman (1956),van Dyke (1957), Hayes & Probstein (1959), and Yakura (1962), to name a few of the earliest works in the hydrodynamics approximation. However, even magnetospheric bow shocks can be studied in the hydrodynamic limit given that the Alfvén and slow rotational magnetosonic waves remain small while the magnetohydrodynamic Mach number is large (Spreiter et al., 1966; Spreiter & Stahara, 1980; Slavin & Holzer, 1981; Verigin et al., 2003; Kotova et al., 2020, 2021). In this work, we are interested in the asymptotic shape of the bow shock far from the obstacle in the astrophysical context, and we highlight the steady-state solution that was explored with direct hydrodynamical simulation in Yalinewich & Sari (2016, hereafter YS16). To find the shape of the bow shock, YS16 propose the thought experiment: replace an object moving through a cold medium by a properly timed series of explosions. Assuming that the resulting blast waves are sandwiched between each other, YS16 get the bow shock shape $z\propto r^{2}$. YS16 was an important motivation for us, although formally speaking we do not use any of their results. Unlike YS16, we derive an exact self-similar asymptotic for the strong bow shock, and we also discuss turbulence that necessarily manifests near the obstacle head.
@@ 2 Theory
One would think that a mathematical description of a bow shock created by a stellar wind is insurmountably difficult. There are at least three surfaces of discontinuity in the flow: two shocks and one tangential discontinuity [^1] The terminology and notations of this section follow Landau & Lifshitz (1959)., see Figure 1. The shapes of these singular surfaces are determined by the flow in between them, and the flow itself depends on the shapes of the singular surfaces. “Mathematically such problems are altogether inaccessible to our present analytical techniques” wrote von Neumann before presenting his exact solution of the strong explosion problem (von Neumann, 1947).
@@ 2.1 Turbulence
Assume that the flow is laminar. Then, in the frame of the star, the flow is steady. In this frame, the cold stellar wind collides with the cold interstellar medium (ISM) wind, see Figure 1. Before colliding, the two winds shock and pressurize. The pressurized ISM and the pressurized stellar gas touch each other along a steady tangential discontinuity surface. The gas velocity normal to the tangential discontinuity is zero, the pressure doesn’t jump across the discontinuity, but the tangential velocity does jump. In three space dimensions, even a finite-Mach tangential discontinuity is unstable. According to (Landau & Lifshitz, 1959, section 84, problem 1), this was shown by Syrovatskii (1954). In an ideal inviscid gas, the instability of the tangential discontinuity manifests as turbulence. The instability and the resulting turbulence are essentially 3D; we will need a 3D simulation to confirm turbulence in Section 3.
@@ 2.2 Asymptotically Self-Similar Bow Shocks
where $\gamma$ is the adiabatic index, assumed to be constant.
@@
As written, Equations (1) – (3) with the boundary conditions (4) – (7) have no meaningful solutions (except $R^{\prime}=\infty$ — a strong planar shock), because the obstacle creating the bow shock is not specified. We will show, however, that at $z\rightarrow+\infty$ a self-similar non-singular asymptotic solution does exist. Namely, for $z\rightarrow+\infty$, one can use the following expansion in small $R^{\prime}$:
@@
This is because, to leading order, $v=1$ at the shock upstream and, to leading order, the Bernoulli equation gives $v={\rm const}$ along a streamline. Thus, asymptotically the bow shock downstream slows down to the ISM velocity. This doesn’t mean that the bow shock becomes “invisible” at great distances from the head. Our bow shock is assumed infinitely strong, so the maximal density at any fixed arbitrarily large $z$ is still a factor of a few (4 for $\gamma=5/3$) greater than the unperturbed ISM density.
Figure 2: The full structure of the bow shock for a rigid obstacle plotted at $t=1,000$. The bow shock wings extend to a maximum of 20 characteristic radii from the symmetry axis.
@@
Equations (15) – (20) are isomorphic to a cylindrically symmetric blast wave; $z$ plays the role of time. An exact solution can immediately be deduced, analogous to the Sedov-von Neumann-Taylor solution (Landau & Lifshitz, 1959, secion 106):
@@
The focal length of the parabola, $f$, is given by the energy deposition rate in the ISM frame [^2] We thank Boaz Katz for pointing this out.. For
@@
Figure 3: The full structure of the bow shock for an obstacle with $v_{\rm wind}=0.1$ plotted at $t=1,000$. The bow shock wings are wider than those of the rigid bow shown in Figure 2 due to inner-outer wind interaction.
@@ 3.2 Results
In Figures 2 and 3 we plot the two edge cases considered in this work. Namely, the full structure of the bow shock for the rigid body and the windiest body, i.e, $v_{\rm wind}=0.1$, respectively. In Figure 2 the density field behind the rigid body is evacuated (i.e., $\rho\ll 1$), a testament to the robustness of the immersed boundary method invoked to model the boundary layer. In Figure 3, the density field only gradually decreases to values $\rho\lesssim 0.1$ at large distances $|z|\geq 150$. This hints at the fact that the windy obstacles approach the asymptotic solutions slower than the perfectly rigid obstacle since the downstream is polluted by turbulent streams of ablated material.
@@ 4 Discussion
D This work has shown that the far asymptotic structure of the bow shock with infinite Mach number is self-similar and and can be described analytically. After derivation of the exact asymptotic solution, we confirm the theory with high-resolution hydrodynamical simulations for a rigid obstacle and a windy obstacle. The rigid obstacle shows near perfect agreement with the theory with a maximal percent difference of 7%. The windy obstacle shows rough agreement with the theory, but this is to be expected since the presence of an inner wind introduces turbulence. As the obstacle gets decelerated, the bow shock shape will transition from a parabola to the canonical Mach cone with opening angle $\theta_{M}=\sin^{-1}(1/M)$. Using Equation (26) in conjunction with the Mach angle, we arrive at
$z_{\rm Mach}$ $=$ $4f(M^{2}-1),$ (33)
$R_{\rm Mach}$ $=$ $4f(M^{2}-1)^{1/2},$ (34)
as the Mach cone transition coordinates. This implies that the assumption that the bow shock remains so strong far from the obstacle is valid in range $R\ll z\ll 4fM^{2}$ for high Mach numbers.
@@
[…]
Moreover, this model may also be applicable to runaway supernova (SN) ejecta blobs that outrun the SN shock front and develop bow shocks like what has been observed in the filamentary structures of the Cassiopeia A SN remnant (Fesen et al., 2006) or the bidirectional ejecta blobs that might develop from explosions of axisymmetric Roche-lobe distorted Type Ib/c SN progenitors (DuPont et al., 2024).
Another potential use case for our model is constraining the sources behind quasi-periodic eruptions (QPEs; Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021; Chakraborty et al., 2021). Some models suggest that QPEs are stars that plunge into active galactic nuclei (AGN; e.g., Lu & Quataert, 2023; Linial & Metzger, 2023; Tagawa & Haiman, 2023). The supersonic star creates a bow shock whose thermal energy is radiation dominated and this model is used to explain the radiative properties of some QPEs. In the future, we plan to utilize our theory to explore the fundamental problem of stars plunging through AGN disks in order to glean some of the physics of the stellar entry and exit. This should provide constraints on the á la mode QPE models proposed.
We thank the referee for providing useful comments which led to the improvement of this work. The authors thank Tamar Faran, Zoltan Haiman, Boaz Katz, and Yuri Levin for useful discussions. MD acknowledges a James Author Fellowship from NYU’s Center for Cosmology and Particle Physics, and thanks the LSST-DA Data Science Fellowship Program, which is funded by LSST-DA, the Brinson Foundation, and the Moore Foundation; his participation in the program has benefited this work.